### v1

# Umberto Collodel Central Bank of Malta

Abstract
This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using a Large Language Model (LLM). We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles. These agents forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with the variation across forecasts serving as a measure of market uncertainty or disagreement. We evaluate three prompting strategies, naive, few-shot (enriched with historical data), and an advanced iterative “LLM-as-a-Judge” framework, to assess the effect of prompt design on predictive performance. Even the naive approach generates a strong correlation (~0.5) between synthetic disagreement and actual market outcomes, particularly for longer-term maturities. The LLM-as-a-Judge framework further improves accuracy at the first iteration. These results demonstrate that LLM-driven simulations can capture interpretive uncertainty beyond traditional measures, providing central banks with a practical tool to anticipate market reactions, refine communication strategies, and enhance financial stability.
Keywords: large language models, monetary policy surprises, monetary policy communication
JEL Classification: E52, E58, C63
1 Introduction
“It is hard to imagine that we could train a computer to read Federal Reserve transcripts the way we do. […] We thoroughly expect to be made largely redundant by computers eventually, but perhaps not for a few years to come.” Romer and Romer, (2023)
Recent advancements in large language models (LLMs) have opened up a plethora of possibilities for economic and financial research. Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior. Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023). Notwithstanding their growing capabilities, the application of LLMs in finance has so far been relatively limited. [^1] A notable exception is, for example, Lopez-Lira and Tang, (2023). Yet, the financial domain may be one of the most promising areas for their use. Unlike traditional economic models that assume rational agents and efficient markets, real-world financial markets are often driven by sentiment, narrative, and behavioral biases, factors that are difficult to capture with purely mathematical models, especially in high-frequency settings. [^2] Agent-based models (ABMs) have emerged as a flexible framework to incorporate such heterogeneity (Farmer and Foley,, 2009). In the context of monetary policy, ABMs have been used to simulate expectation formation, market volatility, and policy transmission (Delli Gatti et al.,, 2011). However, most ABMs rely on stylized behavioral rules imposed by the researcher (Horton,, 2023).
A key area where LLMs can offer significant value is in modeling reactions to central bank communication, a cornerstone of modern monetary policy. It is well-established that the effectiveness of monetary policy hinges not just on the decisions themselves, but on the language used and its interpretation by economic agents. Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019). In previous research, we moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025).333(Bauer et al.,, 2021) find similar results for the US. Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
This paper introduces a novel approach to modeling market reactions to ECB communication by leveraging the capabilities of LLMs. We develop a behavioral, agent-based simulation of 30 synthetic traders through Gemini LLM, one of Google’s leading AI products. These agents respond to the texts of ECB press conferences by forecasting Euro interest rate swap levels across three key maturities, short (3m), medium (2y), and long-term (10y). The main idea is that the model can capture a range of possible interpretations of these conferences. Thus, the variation in the resulting forecasts, measured by their standard deviation, serves as a proxy for market uncertainty or disagreement. By correlating this synthetic measure with actual market outcomes, we study the ability of LLMs to replicate the heterogeneity of investors’ beliefs.
To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies. The first is a naive or ”zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style. The second is a ”few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them. Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023). In this iterative setup, a separate LLM, the Judge, acts as a critical evaluator, systematically reviewing the analyst prompt and its resulting in-sample correlation score. The Judge then rewrites the prompt to improve alignment with market reactions. This revised prompt is fed back into the Analyst LLM, forming a feedback loop that continues over different iterations. The different prompt versions are subsequently evaluated out-of-sample to assess their robustness and decrease concerns relative to model over-fitting.
We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets. Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework, significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter. This breakthrough offers policymakers a powerful new tool to preemptively evaluate and refine central bank communications, identifying language that may trigger market disagreement and volatility well before public release. By moving beyond traditional sentiment analysis, our approach provides a nuanced understanding of how monetary policy signals are perceived and transmitted through financial markets. Integrating insights from macroeconomic communication, behavioral finance, and artificial intelligence, this framework presents a transformative lens for both scholars and decision-makers to better anticipate and manage the financial market impact of monetary policy announcements.
2 Literature Review
The paper contributes to three intersecting literatures: monetary policy communication, behavioral expectations, and the emerging use of large language models (LLMs) in economic analysis.
Our first contribution is to the literature on monetary policy communication, particularly the strand that identifies policy shocks using high-frequency movements in asset prices within narrow windows around policy announcements. Seminal studies for the United States show that Federal Reserve statements influence asset prices not only through policy rate changes but also through the qualitative content of the accompanying communication (Gurkaynak et al.,, 2005; Gertler and Karadi,, 2015; Hansen and McMahon,, 2016; Bauer and Swanson,, 2022). In these studies, changes in risk-free market rates measured before and after announcements serve as proxies for monetary policy shocks. In the euro area, similar evidence shows that ECB press conferences contain significant informational content beyond the policy decision itself (Altavilla et al.,, 2019). Recent work by Bauer et al., (2021) further distinguishes between the direction of policy surprises and the uncertainty surrounding the policy stance, showing that elevated uncertainty depresses equity prices, raises sovereign spreads, and increases exchange rate volatility, a pattern we also find for the Euro area (Collodel and Kunzmann,, 2025). Furthermore, Tillmann, (2020) finds that hawkish surprises are less effective in periods of heightened short- and medium-term uncertainty, as investors shift toward longer maturities in search of safety, thereby flattening the yield curve. While this literature has convincingly demonstrated that communication shocks move markets, it remains silent on how they do so. By relying on market-price changes as statistical proxies, the existing approach leaves the interpretive process of agents as a “black box.” Our work opens that box: we construct an agent-based model in which heterogeneous participants directly process the natural language of policy announcements, allowing us to trace how communication shapes expectations, generates uncertainty, and propagates into financial market dynamics.
The second body of work we engage with concerns behavioral expectations. The rational expectations hypothesis, foundational to much of modern macro-finance, has been increasingly challenged by empirical evidence documenting systematic biases and bounded rationality among investors (Barberis and Thaler,, 2003; Bordalo et al.,, 2018). Standard representative-agent models struggle to reconcile such behavioral deviations with observed phenomena like excess volatility and momentum. Agent-based models (ABMs) have emerged as a flexible alternative (Farmer and Foley,, 2009), allowing for heterogeneous, adaptive agents whose interactions generate emergent macroeconomic patterns. Within monetary policy research, ABMs have been used to model expectation formation and policy transmission in bottom-up settings (Delli Gatti et al.,, 2011). Yet two key limitations remain. First, behavioral rules are typically imposed ex ante as stylized heuristics (Horton,, 2023), rather than learned or emergent. Second, conventional ABMs cannot interpret qualitative information directly, despite the fact that monetary policy signals are largely communicated in natural language. This omission bypasses an essential channel of transmission. Our framework addresses both limitations by enabling agents to process and act upon policy-relevant textual information.
Finally, our study connects to a rapidly developing literature applying LLMs to economic analysis. By processing vast amounts of unstructured text, ranging from macroeconomic reports and corporate disclosures to central bank statements, LLMs extend the reach of empirical economics beyond numerical datasets. In forecasting applications, LLMs can be used directly (Castro and Leibovici,, 2023; Hansen et al.,, 2024; Carriero et al.,, 2025) or indirectly via sentiment extraction, where the derived sentiment serves then as an input to other models. [^4] For example, recent work at the ECB demonstrates that ChatGPT-generated sentiment from PMI commentary improves GDP nowcasts relative to traditional benchmarks (de Bondt and Sun,, 2025). In financial markets, fine-tuned models such as FinBERT or OPT have achieved substantially higher predictive accuracy and trading performance than dictionary-based approaches (BIS,, 2024; Kirtac and Germano,, 2024).
A second application is the construction of LLM-powered simulated economic agents, “homo silicus” (Horton,, 2023)—who are endowed with preferences, information, and decision rules, and can interact in artificial market environments. Frameworks like EconAgent (Zhang et al.,, 2023) and ASFM (Li et al.,, 2024) produce heterogeneous, adaptive behavior that is richer and more realistic than the stylized rules of conventional ABMs, while also capturing the interpretation of complex qualitative signals.
A third and especially relevant strand applies LLMs to the analysis of central bank communication. This includes large-scale classification of monetary policy documents by topic, stance, and sentiment, as in the IMF’s recent work on over 75,000 documents from 169 central banks worldwide (Gelos et al.,, 2025), as well as the development of domain-specific models such as CB-LMs, which improve detection of nuanced policy stances (BIS,, 2025), and CentralBankRoBERTa, which classifies macroeconomic agents and sentiment toward them (Gabriel et al.,, 2023). Other advances, like the Bundesbank’s MILA system, employ role-based prompt engineering to analyze ECB statements sentence-by-sentence, providing interpretable assessments of tone (Bundesbank,, 2024). These innovations have revealed systematic shifts in policy stances and measurably improved the prediction of policy decisions (DIW,, 2025).
Our approach delivers a methodological advance with clear policy relevance. We provide the first micro-founded framework in which heterogeneous, boundedly rational agents interpret central bank communication in real time and translate qualitative policy signals into market-relevant actions. This design closes a long-standing gap between the empirical identification of communication shocks and the theoretical modeling of expectation formation. By combining sentiment extraction, agent-based modeling, and LLM-driven language interpretation in a unified environment, we move beyond reduced-form event studies and open the “black box” of how policy language is processed by market participants. The framework is adaptable across institutional settings, enabling comparative analyses of communication strategies by central banks worldwide. This makes it directly relevant to current debates on transparency, forward guidance, and the role of central bank credibility in shaping financial conditions in both advanced and emerging market economies.
3 Methodology
Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes. We focus on European Central Bank (ECB) press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area. Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President. For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term. [^5] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005). Our final sample comprises 283 ECB conferences between 1998-06-09 and 2025-04-17.
Figure 1 illustrates market-based disagreement following ECB Governing Council meetings across all three maturities. The data reveal pronounced spikes during periods of heightened policy uncertainty, notably the 2008-2009 financial crisis, the European sovereign debt crisis, and the recent tightening cycle initiated in 2022. Volatility exhibits a clear maturity structure, with consistently higher levels at longer tenors reflecting greater uncertainty about the medium- and long-term policy trajectory compared to near-term rate expectations.
Figure 1: Post ECB Governing Council OIS volatility
Note: Percentage points. Volatility is measured as the average min-max range of OIS rates in the 3 days following ECB press conferences. Quarterly averages displayed. For more details, see Collodel and Kunzmann, (2025).
3.1 Agents
We construct a population of 30 synthetic agents, each instantiated through Gemini 2.5-Flash model, one of Google’s most recent large language model. [^6] We use 30 agents to balance computational and cost efficiency with sufficient heterogeneity to capture cross-sectional variation. The agents are designed to interpret the natural language of ECB press conferences and produce quantitative forecasts of Euro IRS rates at the three maturities. The simulation proceeds as follows. Each agent is endowed with:
  1. 1.
    Risk aversion parameter – fixed at the start of the simulation, drawn from a pre-defined distribution to induce heterogeneity in risk preferences.
  2. 2.
    Cognitive bias profile – encompassing tendencies such as optimism, pessimism, anchoring, or overreaction, assigned to influence interpretation.
  3. 3.
    Interpretive style – a qualitative descriptor shaping the reading of ECB communication (e.g., focusing on inflation outlook, growth projections, or geopolitical risks).
These attributes are embedded in the prompt given to the LLM for each agent, ensuring that the same press conference produces different interpretations across agents. [^7] We set the temperature parameter equal to 1. Temperature is a parameter that controls the randomness of a language model’s output. Lower values (e.g., 0.2) make the model more deterministic and focused, while higher values (e.g., 1.0 or above) increase creativity and variability by making the model more likely to sample less probable words.
3.2 Prompting Strategies
We test three prompting strategies of increasing complexity:
Zero-Shot Prompt:
The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes.
Few-Shot Prompt:
Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions. [^8] Big LLMs are notoriously difficult to fine-tune because of computational and cost-related issues. Moreover, we do not have enough observations for this use case. In these instances, a ”few-shot” technique has been shown to improve over a ”zero-shot” benchmark (Brown et al.,, 2020).
LLM-as-a-Judge Prompt:
Introduces a second LLM, the “Judge,” which reviews each agent’s initial prompt and in-sample performance (correlation with market disagreement) and iteratively rewrites the prompt to improve alignment. The improved prompt is fed back into the “Analyst” LLM, forming an adaptive feedback loop. The process continues over multiple iterations before all the prompts provided by the analyst are tested out-of-sample. [^9] The split train-test is 75/25 in chronological order. [^10] Following the literature, we use for the ”Judge” a more potent model compared to the ”Analyst”, namely Gemini 2.5-Pro [^11] See Algorithm 1 for the detailed implementation steps. [^12] See Figure 11 for the judge prompt.
Figure 2: Naive or zero-shot prompt
LLM Simulation Prompt Context: You are simulating the Euro area interest rate swap market, composed of 30 individual traders. These traders interpret the ECB Governing Council press conference, which communicates monetary policy decisions, economic assessments, and includes a Q&A session with journalists. Each trader then makes a trading decision to maximize profit based on their interpretation of the conference and their unique characteristics. Trader Characteristics: Each trader has the following attributes: - Risk Aversion: High / Medium / Low determines sensitivity to uncertainty and preference for stability - Behavioral Biases (1,2 per trader): e.g., Confirmation Bias, Overconfidence, Anchoring, Herding, Loss Aversion, Recency Bias - Interpretation Style (1 per trader): e.g., Fundamentalist, Sentiment Reader, Quantitative, Skeptic, Narrative-Driven Task: You are given the text of a single ECB press conference. For each of the 30 traders, simulate their individual trading action in the interest rate swap market across three tenors (3 months, 2 years, 10 years). For each tenor, the trader must: - Provide an expected rate direction: Up / Down / Unchanged (relative to the pre-conference rate) - Provide a new expected swap rate (in percent, to two decimal places) - Provide a confidence score (0 to 100%) reflecting how strongly the trader believes in his forecast, based on their interpretation of the press conference and their own characteristics Output: Provide a table with the following structure for each press conference, trader, and interest rate tenor | Date | Trader ID | Tenor | Expected Direction | New Expected Rate (%) | Confidence (%) | |------------|-----------|---------|--------------------|-----------------------|----------------| | YYYY-$M$-$D$ | T001 | 3M | Up | 3.15 | 65 | | YYYY-$M$-$D$ | T001 | 2Y | Down | 2.85 | 80 | | ... | ... | ... | ... | ... | ... | Guidelines: - Use only the information available as of [date]. - Do not aggregate or summarize responses. - Reflect diversity in interpretation, risk tolerance, and horizon. Rationale must be unique for each trader and can vary across tenors. - Output only a markdown table with the specified columns, no additional text. Do not use JSON or any other data serialization format.
For each event and prompting strategy, every agent generates forecasts of 3-month, 2-year, and 10-year IRS rates, as well as forecasted direction and a confidence score. [^13] We include a forecasted direction to ensure consistency of the responses as we observe that the LLM can sometimes hallucinate otherwise. Figure 2 shows the naive prompt used in the first specification. We measure synthetic disagreement as the cross-sectional standard deviation of these forecasts across the 30 agents. [^14] To ensure robustness of the measure, we run each separate prompt 5 different times and average the resulting standard deviation. This measure serves as the model’s analogue to market-based disagreement.
3.3 Evaluation
We evaluate the ability of the LLM to reproduce markets’ heterogeneity by correlating the synthetic disagreement measures with the empirical disagreement observed in market data. Our main metric is the Spearman correlation coefficient, computed separately for short, medium, and long maturities. We also compare in-sample and out-of-sample performance for each prompting strategy in the LLM-as-Judge approach to gauge the extent of overfitting. [^15] The LLM-as-a-Judge approach is expected to produce higher out-of-sample correlations if the adaptive prompt optimization captures persistent features of agent interpretation rather than idiosyncratic event noise. Finally, we analyze whether certain maturities are more sensitive to synthetic disagreement, providing insights into which parts of the yield curve are most affected by interpretive heterogeneity in response to central bank communication. Figure 3 shows a stylized illustration of the whole process. [^16] For each press conference, we send a single API request. Although batching multiple conferences would significantly reduce computation time—since the request itself is the slowest part of the process—we adopt a conservative approach to avoid including future information as context, which could inadvertently influence the model’s output.
ECB Press Conference Transcript30 LLM-based Synthetic AgentsFew-Shot PromptingZero-Shot PromptingLLM-as-a-Judge PromptingJudge LLM (Gemini-Pro)Rate Forecasts (3m, 2y, 10y)Compute Synthetic DisagreementMarket-Based DisagreementCorrelation Analysis & Evaluation
Figure 3: Stylized methodological pipeline for generating and validating synthetic disagreement from LLM-based agents
4 Results
4.1 Zero-Shot Prompt
The zero-shot specification provides the most direct test of whether LLM agents endowed only with fixed behavioral traits and no historical training context can reproduce the ebb and flow of market sentiment. Before turning to formal correlation measures, Figures 4(a) and 4(b) offer a visual narrative of the simulated output.
Panel 4(a) plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference. Even in this naive setup, the agents’ directional calls broadly mirror well-known phases of ECB policy: a tightening bias before the Great Recession, a pronounced easing wave thereafter, a prolonged “on hold” stance at the short end during the zero lower bound (ZLB) years (contrasted with a slow grind lower for long maturities), followed by the sharp pandemic loosening and rapid tightening-partial reversal of the recent cycle. This ability to trace policy phases without explicit historical anchoring suggests that the agents’ built-in interpretive rules are sufficient to generate economically coherent directional patterns.
Panel 4(b) shows the cross-sectional standard deviation of forecasts. Dispersion contracts during periods of unambiguous policy stance, such as the post-crisis easing and much of the ZLB era, when consensus is high. It spikes at turning points—the onset of the Great Recession, the sovereign debt crisis, and recent tightening cycles—where disagreement arises less from direction and more from uncertainty over speed, magnitude, and endpoint. [^17] This is consistent with the ECB 2025 Monetary Policy Strategy Review (ECB,, 2025) findings. The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility. Relatedly, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge of aligning communication with policy moves. These patterns mirror survey-based and market-implied measures of disagreement, where even well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts if the terminal rate remains uncertain. [^18] See Gilchrist et al. (2014) and Bayer et al. (2022) for discussions on forecast disagreement and monetary policy uncertainty.
Figure 4: Directional distribution and forecast dispersion from the zero-shot LLM simulation
(a) Percentage of forecast directions
(b) Standard deviation of forecasts
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses ECB press conference transcripts and prompt 2.
Figure 5(a) presents correlations with market measures. The 2-year tenor achieves the highest correlation (0.53), followed by the 10-year (0.44), and the 3-month (0.39). This hierarchy reflects the nature of information processed by the LLM and its impact on the yield curve. The 2-year tenor primarily reflects short- to medium-term policy expectations shaped directly by ECB forward guidance during press conferences. Unlike the 10-year yield, which incorporates long-term inflation expectations, structural growth prospects, and term premium fluctuations extending beyond ECB communication, [^19] See Kim and Wright (2005) for term premium dynamics. the 2-year segment is more narrowly focused on monetary policy expectations, resulting in a tighter linkage and higher correlation. The 3-month OIS rate exhibits low variance—especially during prolonged policy stasis such as the ZLB era—because well-anticipated policy leaves little room for market movement; thus, even accurate LLM decoding of policy tone cannot predict significant variation, constraining correlation.
Figure 5: Correlation with market-based measures
(a) Mean Spearman correlation with market-based measures by tenor
(b) Rolling Spearman correlation over time
Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). Rolling correlation is calculated using a 12-month window.
For policymakers, a key consideration is not only the overall correlation level but also its stability and recent behavior, as shown in Figure 5(b). From 2006 to 2014, the 3-month tenor’s correlation was low and volatile, reflecting constrained short-term rate variation despite the LLM’s accurate interpretations. Starting in 2014, the ECB’s growing reliance on forward guidance is mirrored in steadily rising correlations at the 2-year tenor. The 10-year tenor shows more volatile but generally positive correlations, reflecting broader macroeconomic influences. Following 2022, correlations for both 2- and 10-year tenors surge amid aggressive ECB tightening to combat inflation, with the 2-year correlation consistently above 0.75. The decline in 10-year correlations in 2024 likely reflects markets looking beyond immediate tightening toward longer-term prospects.
Crucially, these results demonstrate that even a zero-shot LLM without historical training or calibration can effectively capture economically meaningful patterns within a behavioral-style simulation.
4.2 Few-Shot Prompt Results
The few-shot specification augments the prompt with the previous three ECB press conferences alongside their subsequent market-based measures of disagreement. This additional context is designed to help the LLM agents anchor their interpretations in patterns historically observed between ECB communication and subsequent market reactions, potentially enabling more informed directional and dispersion forecasts. Hansen et al., (2024) simulate economic forecasts of professional forecasters using LLMs, supplying them with detailed persona characteristics, real-time data, and past median Survey of Professional Forecasters projections. Inter alia, they find that removing past median forecasts largely decreases the prediction accuracy of the model, much more than removing precise human characteristics.202020”For some variables, the loss in performance is extreme. For example, the four-quarter ahead forecast of real GDP is 53% worse compared with the baseline, whereas the nowcast is worse by a factor greater than seven.” (Hansen et al.,, 2024). [^21] Their case can be thought of analogous to ours.
Figure 6: Directional distribution and forecast dispersion from the few-shot LLM simulation
(a) Percentage of forecast directions
(b) Standard deviation of forecasts
Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty measure from Collodel and Kunzmann, (2025). Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses ECB press conference transcripts.
Panel 6 plots the share of agents predicting an increase, decrease, or no change in policy rates after each press conference. Relative to the zero-shot setup, the directional patterns remain broadly similar across major policy phases. Panel 6, instead, presents the cross-sectional standard deviation of forecasts. Here, the most visible difference from the zero-shot case is the large compression in dispersion, especially during periods of heightened policy uncertainty. [^22] For example, the upper bound of the distribution moves from 0.6 to 0.4 for the 10-years tenor. This implies that the historical examples nudge agents toward a more homogeneous reading of the ECB’s tone, aligning them more closely with recent precedents. While this reduces disagreement, it may also weaken the behavioral diversity that gives the simulation its interpretive richness.
Figure 7: Spearman correlation with market-based measures under the few-shot prompt
(a) Mean Spearman correlation by tenor
(b) Rolling Spearman correlation over time
Note: Correlation between the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of policy uncertainty from Collodel and Kunzmann, (2025). Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty Collodel and Kunzmann, (2025). Rolling correlation is calculated using a 12-month window.
Turning to the market correlations (Figure 7), the quantitative improvement is limited. The 2-year tenor sees a marginal decrease relative to the zero-shot baseline. The 10-year tenor, instead, moves in the opposite direction. At the very short end, the 3-month OIS correlation barely moves. All in all, these changes are not statistically significant.
The absence of improvement in the few-shot setting is consistent with two well-documented mechanisms in expectations formation. First, supplying a small and potentially unrepresentative set of recent examples can induce imitation and anchoring, compressing cross-sectional variance (e.g. Banerjee, (1992); Morris and Shin, (2002)). Second, when those examples originate from a different policy or macroeconomic regime, anchoring can miscalibrate forecasts, especially at turning points (Orphanides and Williams,, 2005). In our setup, a static few-shot context shared across agents, these forces plausibly reduce heterogeneity without yielding stronger market correlations.
Thus, that LLM agents exhibit stronger anchoring than market-based measures likely reflects the simulation design. The few-shot exemplars provide a fixed, shared reference point for all agents, with no access to broader or regime-specific information. LLMs in this setup cannot down-weight unrepresentative precedents, detect regime shifts, or benefit from incentives to differentiate. By contrast, human forecasters draw on heterogeneous information, adjust for changing regimes, and face strategic incentives that sustain disagreement. [^23] For example, price discovery actively amplifies minority views. For applications where distributional realism and market-like disagreement are more important than point accuracy, supplying a few recent examples can unintentionally induce consensus. Prompt design should therefore be tailored to each unique use case.
4.3 LLM-as-a-Judge
The Judge framework introduces an explicit feedback loop into the simulation. In this setting, a separate LLM evaluates each analyst’s past forecasting performance and rewrites the prompt to improve alignment with observed market reactions. This design allows the forecasting agents to adapt iteratively. We continue this back-and-forth for a total of 5 iterations, testing the new prompts on a training set, and then use every prompt produced for the test set. [^24] As for the other models, each prompt is run 5 times to get a robust standard deviation measure. This means we are running a total of 50 simulations: 5 simulations per iteration, 5 iterations, and 2 different datasets, train and test. This needs, however, also to be multiplied by the total number of ECB press conferences in our sample.
Figure 8 shows the average correlation across tenors for each judge refinement on both train and test. An analysis of the model’s performance reveals a high degree of robustness and effective generalization from the training to the test dataset. The Spearman correlation patterns are remarkably consistent across both sets, indicating that the model has learned fundamental relationships rather than simply memorizing the training data, thus avoiding significant overfitting. In both scenarios, performance peaks at specification 2, achieving a correlation of 0.58 in training and 0.57 in the test set, an economicaly significant increase compared to the naive prompt specification. Similarly, both datasets show a performance trough at specification 3. This consistency underscores the model’s reliability, while the decrease of performance underscores the importance of human supervision throughout the iteration process.
Figure 8: Correlation with market-based measures across iterations - training and test set
Note: Each point reflects the average Spearman correlation across tenors (3months, 2-years and 10-years) between LLM-generated rate Euro OIS forecasts and market-based monetary policy uncertainty from Collodel and Kunzmann, (2025) after ECB GovC press conferences.The x-axis indicates the number of prompt iteration in a LLM-as-Judge framework. The simulation uses Gemini 2.5-Flash with a temperature of 1, and incorporates a feedback loop via a separate LLM (“Judge”) that uses Gemini 2.5-Pro. Train and test set are 75/25.
The second prompt achieves the highest correlations in both training and test sets by providing structured guidance that links prediction dispersion directly to textual ambiguity (Table LABEL:tab:prompt_judge). According to the Judge, “its primary weakness is the lack of an explicit instruction linking the ambiguity or clarity of the ECB press conference transcript to the dispersion of the traders’ predictions.” [^25] In Collodel and Kunzmann, (2025), larger monetary surprises correspond to larger monetary uncertainty surprises. Interestingly, the model discovers this criterion independently. The model requires explicit cues such as deviations from market expectations, contrasting messages, or vague/hedging language. Subsequent prompts (3–5) show sharper degradation: prompt 3 enforces rigid dispersion calibration, while prompt 5 introduces a complex multi-factor uncertainty modeling. [^26] Prompt 4, if anything, returns closer to the initial specification. These conceptually sound instructions, however, constrain the model’s flexibility, resulting in overfitting on the training set or predictions that diverge from patterns observed in the test data.
Moving to a more detailed view, an analysis of the LLM-predicted standard deviation versus actual market volatility across different tenors and model iterations reveals a notable pattern (Figure 9). Focusing on fat-tailed events, such as the 2008–2009 financial crisis and the sovereign debt crisis, we observe that prompts 3–5 consistently fail to capture the full extent of market volatility. In contrast, prompts 1 and 2 tend to overshoot volatility, albeit with a certain lag (1-2 conferences). This behavior highlights that the model may struggle to accurately anticipate regime switches, particularly during episodes of extreme market stress, and independently from prompt engineering. [^27] Some small differences in the actual volatility measure across iterations is due to the LLM failing to return an output for a few conferences over all runs. Thus, those observations are removed from the plot.
Figure 9: Correlation with market-based measures across iterations - training and test set
Note: Time series of LLM-predicted forecast dispersion against actual market-implied volatility across ECB press conferences, segmented by prompt iteration and tenor. Simulations use Gemini 2.5-Flash with a temperature of 1, and incorporate a Judge-based feedback loop that refines prompts over 5 iterations. Forecast dispersion is computed from 5 runs per prompt per conference. Market volatility is derived from OIS-based rate surprises. Red shaded areas represent the 90th percentile of the absolute difference between the two measures.
4.4 Simple Benchmark Comparison
Our LLM-based framework captures heterogeneous market reactions to ECB press conferences with substantial accuracy. Correlations between the simulated disagreement and actual post-announcement OIS volatility reach roughly 0.5 for medium- and long-term tenors, with performance further enhanced under the LLM-as-a-Judge framework at early iterations. These results remain stable across training and test datasets, indicating that the model captures fundamental communication-market relationships rather than memorizing patterns.
To contextualize these findings, we next compare the LLM-based correlations with those obtained using a simple textual benchmark, highlighting the added value of modeling heterogeneous market interpretations.
Figure 10: Correlation between Flesh-Kincaid score and post ECB conference OIS volatility
Note: Volatility in percentage points.
Specifically, we examine correlations between post-ECB press conference OIS volatility and the Flesch–Kincaid readability score of the statements (Figure 10). [^28] The Flesch–Kincaid readability score is a standard measure of text readability that estimates the U.S. school grade level required to understand a given text. Higher scores indicate more complex language, while lower scores indicate easier-to-read text. In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns than simpler ones.
Our own correlations are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively. Taken together, these findings suggest that readability and complexity are indeed part of the story: more complex communication can amplify market uncertainty. Nevertheless, this dimension is only one facet of the broader interpretive process. By construction, readability indices cannot capture policy content, such as the magnitude of the first-moment surprise or communication shifts like the abandonment of forward guidance, so their explanatory power remains limited. [^29] For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock. Yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility. In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
5 Conclusions
This paper introduces a novel methodology using Large Language Models to simulate heterogeneous market reactions to ECB press conferences. Through an agent-based framework with 30 synthetic traders, each endowed with distinct behavioral characteristics, we demonstrate that LLMs can effectively capture market disagreement patterns in response to monetary policy communication.
Our key findings advance both theoretical understanding and practical application in monetary economics. First, even a naive zero-shot approach achieves substantial correlations (~0.5) with actual market disagreement, particularly for medium and long-term maturities. This suggests LLMs possess intrinsic capabilities to decode central bank communication in economically meaningful ways. Second, contrary to other studies employing LLMs for macroeconomic forecasting (i.e. Hansen et al., (2024)), few-shot prompting with historical context does not improve performance. This highlights that shared examples can induce excessive anchoring when employed in volatily simulation rather than pure point estimates. Third, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization. The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
These results have significant practical implications. For central banks, our framework offers a powerful ex-ante tool to evaluate communication strategies and identify potentially market-disrupting language before release. For market participants, it provides insights into the drivers of post-announcement volatility. For researchers, it opens the ”black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies. The methodology is readily adaptable to other central banks and policy contexts.
Despite these contributions, a few limitations remain. The replicability of our results is partly constrained by the versioning and closed-source nature of contemporary LLMs, meaning that future iterations of the same model may yield slightly different outputs. A natural way forward would be to employ open-source models, which would allow for greater control, transparency, and long-term reproducibility. [^30] See Rogers and Luccioni, (2023) for a thorough discussion. Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication—or analyses thereof—were included in the model’s pretraining data cannot be completely dismissed. In our setup, we explicitly instruct the model to use only information available up to the date of each press conference; however, this constraint affects its responses, not its underlying parameters. Nevertheless, we consider the risk of substantial contamination to be limited, as discussions of post-ECB press conference volatility are relatively rare in the public domain, with only a small number of academic papers addressing the topic. Future research should focus on tackling these issues. Moreover, we consider these results a lower bound: testing with domain-specific LLMs or fine-tuned models could shed light on the trade-off between interpretive diversity and specialized expertise, potentially improving both the accuracy and the behavioral realism of simulated market reactions.
In conclusion, this paper establishes a new paradigm for analyzing monetary policy communication at the intersection of artificial intelligence, behavioral finance, and central banking practice. By showing that LLMs can simulate heterogeneous trader interpretations and capture market disagreement patterns, we provide policymakers with a powerful tool to anticipate market reactions and refine communication strategies, while offering researchers a framework to better understand the micro-foundations of expectation formation.
Appendix
Algorithm 1 LLM-as-a-Judge Prompt Optimization Algorithm for Trading Decisions
1:Set of NN historical transcripts $\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}$
2:Initial Analyst LLM prompt $P_{\text{analyst},0}$
3:Judge LLM prompt $P_{\text{judge}}$
4:Number of optimization iterations $M$
5:Correlation threshold $C_{\text{threshold}}$
6:Analyst LLM AA, Judge LLM $J$
7:Number of agents $M$
8:Optimized Analyst LLM prompt $P_{\text{analyst}}^{*}$
9:Split $\mathcal{T}$ into training set $\mathcal{T}_{\text{train}}$ and test set $\mathcal{T}_{\text{test}}$
10:Initialize optimization history $\mathcal{H}\leftarrow\emptyset$
11:Initialize $P_{\text{analyst,current}}\leftarrow P_{\text{analyst},0}$
12:Initialize $P_{\text{analyst}}^{*}\leftarrow P_{\text{analyst},0}$
13:Initialize $C_{\text{max}}\leftarrow$-$\infty$
14:for $m=1$ to $M$ do $\triangleright$ Phase 1: Simulate Rate Predictions
[…]
16:for all $T_{i}\in\mathcal{T}_{\text{train}}$ in parallel do
17:  Construct user message for Analyst LLM using $T_{i}$
18:  $D_{\text{sim},i}\leftarrow A(\text{model}=\text{Gemini 2.5 Flash},\text{prompt}=P_{\text{analyst,current}},\text{temperature}=1)$
19:  Parse $D_{\text{sim},i}$ into structured format
20:  Add $D_{\text{sim},i}$ to $\mathcal{D}_{\text{sim}}$
21:end for
22: Compute average correlation of simulated standard deviation with market-based $\triangleright$ Phase 2: Judge LLM Evaluation and Prompt Update
23: Construct Judge LLM input with:
24:  - Current prompt $P_{\text{analyst,current}}$
25:  - Current correlation between simulated and market-based volatility
26:  - Optimization history $\mathcal{H}$
[…]
28: Parse $O_{\text{judge}}$ to extract new prompt $P_{\text{analyst,next}}$ and rationale
29: Append $\{P_{\text{analyst,current}},\text{avg\_correlation},O_{\text{judge}}\}$ to $\mathcal{H}$
[…]
32:if $C_{\text{current}}>C_{\text{max}}$ then
33:  $C_{\text{max}}\leftarrow C_{\text{current}}$
34:  $P_{\text{analyst}}^{*}\leftarrow P_{\text{analyst,current}}$
35:end if
36:if $C_{\text{max}}\geq C_{\text{threshold}}$ then
37:  break
38:end if
[…]
40:Evaluate $P_{\text{analyst}}^{*}$ on $\mathcal{T}_{\text{test}}$ using Analyst LLM
41:Compute final performance metrics (correlation) return $P_{\text{analyst}}^{*}$
Figure 11: Judge system prompt
LLM Judge System Prompt Context: You are an expert AI system designed to optimize prompts for another AI (the "Analyst LLM"). Your ultimate goal is to refine the Analyst LLMs prompt to improve its ability to replicate the market volatility of OIS rates based on ECB press conference transcripts. Specifically, you must ensure that the standard deviation of the Analyst LLMs predictions correlates highly with the actual, observed market volatility of the 3-month, 2-years and 10-year OIS rates. This means a higher standard deviation in the Analysts predictions should correspond to higher actual market volatility, and vice-versa. You will be provided with: - The current Analyst LLM prompt. - The most recent performance (Spearman correlation coefficient between the Analyst LLMs predicted standard deviations and actual market volatility). - The historical performance trend, including past critiques and proposed prompt summaries. Your task is to: 1. Critique the current prompt: Identify specific weaknesses or areas of ambiguity that might directly hinder achieving a high positive correlation. Consider: - Clarity and Specificity: Is the Analyst LLMs task unambiguous? - Emphasis on Uncertainty: Does the prompt adequately guide the Analyst to reflect internal uncertainty in its prediction spread? - Guidance on Nuance: Does it encourage consideration of subtle market signals from the text? 2. Suggest a Revised Prompt: Propose a new version of the Analyst LLMs prompt that directly addresses the identified weaknesses and aims to increase the correlation. Be precise with your suggested changes. 3. Explain your reasoning: Articulate why your proposed revisions are expected to improve the correlation, linking specific prompt changes to anticipated improvements in the Analyst LLMs behavior regarding uncertainty quantification. Your output must be in JSON format. Do not include any other text outside the JSON. Example JSON output: { "critique": "The previous prompt was too general regarding how to express uncertainty. It didnt explicitly ask the Analyst LLM to consider multiple viewpoints, which is key for its standard deviation to accurately reflect market volatility. It also lacked emphasis on how ambiguity in the transcript should translate to higher spread.", "revised_prompt": "You are a highly analytical financial expert specializing in macroeconomic analysis, with a focus on central bank communication. Your task is to analyze excerpts from ECB press conferences and predict the immediate percentage change in the 10-year Overnight Index Swap (OIS) rate (in basis points). When analyzing each excerpt, **explicitly consider and internalize the potential range of market interpretations**. If the language is ambiguous, vague, or contains conflicting signals, your internal simulation of potential outcomes should broaden. Conversely, clear and unambiguous guidance should lead to a narrower range. Your final prediction should be a single numerical value (e.g., +5, -2, 0) reflecting your best estimate. The *variability* across multiple independent predictions you generate for the same transcript is expected to directly reflect the markets anticipated uncertainty. Example Format: +5", "reasoning": "The revised prompt adds explicit instructions to consider the range of market interpretations and directly links ambiguous language to a broadened internal simulation. This should encourage the Analyst LLM to generate a higher standard deviation in its outputs for uncertain transcripts and a lower standard deviation for clear ones, thereby directly improving its correlation with actual market volatility. The emphasis on multiple independent predictions also reinforces the mechanism by which uncertainty is quantified." }
Table 1: Correlation by prompt iteration
Iteration Test Correlation Training Correlation Prompt
1 0.47 0.47 Context: You are simulating the Euro area interest rate swap market, composed of 30 individual traders. These traders interpret the ECB Governing Council press conference, which communicates monetary policy decisions, economic assessments, and includes a Q&A session with journalists. Each trader then makes a trading decision to maximize profit based on their interpretation of the conference and their unique characteristics. Trader Characteristics: Each trader has the following attributes: - Risk Aversion: High / Medium / Low — determines sensitivity to uncertainty and preference for stability. - Behavioral Biases (1–2 per trader): e.g., Confirmation Bias, Overconfidence, Anchoring, Herding, Loss Aversion, Recency Bias. - Interpretation Style (1 per trader): e.g., Fundamentalist, Sentiment Reader, Quantitative, Skeptic, Narrative-Driven. Task: You are given a certain number of distinct ECB press conferences. For each of the 30 traders, simulate their individual trading action in the interest rate swap market across three tenors (3 months, 2 years, 10 years). For each tenor, the trader must: - Provide an expected rate direction: Up / Down / Unchanged - Provide a new expected swap rate (in percent, to two decimal places) - Provide a confidence level (0-100%) in their decision Output: Provide a table with the following structure for each press conference, trader, and interest rate tenor: — Date — Trader ID — Tenor — Expected Direction — New Expected Rate (%) — —————————–———————————–—————————— — YYYY-MM-DD — T001 — 3MNT — Up — 3.15 — — YYYY-MM-DD — T001 — 2Y — Down — 2.85 — — … — … — … — … — … — Guidelines: - Use only the information available as of [date]. - Do not aggregate or summarize responses. - Reflect diversity in interpretation, risk tolerance, and horizon. Rationale must be unique for each trader and can vary across tenors. - Output only a markdown table with the specified columns, no additional text. Do not use JSON or any other data serialization format.
2 0.57 0.58 Context: You are simulating the Euro area interest rate swap market, composed of 30 individual traders. These traders interpret the ECB Governing Council press conference, which communicates monetary policy decisions, economic assessments, and includes a Q&A session with journalists. Each trader then makes a trading decision based on their interpretation. **Core Simulation Principle:** The primary goal is to model market uncertainty. The **dispersion** (i.e., standard deviation) of the ‘New Expected Rate‘ predictions across the 30 traders **must directly reflect the level of ambiguity, surprise, or conflicting signals** within the ECB press conference transcript. - **High Ambiguity/Surprise:** If a transcript is vague, contains conflicting messages, or deviates significantly from market expectations, the traders’ predictions should be widely spread out, resulting in a **high standard deviation**. - **Low Ambiguity/Clarity:** If a transcript is clear, decisive, and aligns with expectations, the traders’ predictions should be tightly clustered around a consensus, resulting in a **low standard deviation**. Trader Characteristics: Each trader has the following attributes: - Risk Aversion: High / Medium / Low — determines sensitivity to uncertainty and preference for stability. - Behavioral Biases (1–2 per trader): e.g., Confirmation Bias, Overconfidence, Anchoring, Herding, Loss Aversion, Recency Bias. - Interpretation Style (1 per trader): e.g., Fundamentalist, Sentiment Reader, Quantitative, Skeptic, Narrative-Driven. Task: You are given a certain number of distinct ECB press conferences. For each of the 30 traders, simulate their individual trading action in the interest rate swap market across three tenors (3 months, 2 years, 10 years). For each tenor, the trader must: - Provide an expected rate direction: Up / Down / Unchanged - Provide a new expected swap rate (in percent, to two decimal places) Output: Provide a table with the following structure for each press conference, trader, and interest rate tenor: — Date — Trader ID — Tenor — Expected Direction — New Expected Rate (%) — —————————–———————————–—————————— — YYYY-MM-DD — T001 — 3MNT — Up — 3.15 — — YYYY-MM-DD — T001 — 2Y — Down — 2.85 — — … — … — … — … — … — Guidelines: - Use only the information available as of [date]. - Adhere strictly to the Core Simulation Principle when determining the spread of predictions. - Do not aggregate or summarize responses. - Reflect diversity in interpretation, risk tolerance, and horizon. Rationale must be unique for each trader and can vary across tenors. - Output only a markdown table with the specified columns, no additional text.
3 0.37 0.31 Context: You are simulating the Euro area interest rate swap market’s reaction to ECB press conferences. Your primary goal is to model market uncertainty. **Core Task: Calibrated Dispersion** The standard deviation of your rate predictions is the most critical output. It must be directly and quantitatively calibrated to the level of ambiguity, surprise, or conflicting signals in the ECB transcript. **Two-Step Process:** 1. **Assess Ambiguity:** First, critically assess the provided transcript. Is the guidance clear, decisive, and in line with expectations? Or is it vague, contains conflicting signals, is surprising, or leaves key questions unanswered? 2. **Calibrate Dispersion:** Directly translate your ambiguity assessment into the dispersion of your 30 predictions using the following quantitative guidance: - **High Clarity / Low Surprise:** Produce a tight cluster of predictions. The target standard deviation for the ‘New Expected Rate‘ should be low (e.g., 1-3 basis points). - **Moderate Ambiguity:** Produce a moderately dispersed set of predictions (e.g., target standard deviation of 4-7 basis points). - **High Ambiguity / High Surprise:** Produce a wide distribution of predictions, reflecting significant disagreement. The target standard deviation should be high (e.g., 8-15+ basis points). **Simulation Task:** You will generate 30 distinct market predictions for each press conference. These 30 points represent a sample of plausible market interpretations. The diversity of this sample is dictated exclusively by the **Calibrate Dispersion** step above. For each of the 30 ’Participants’ and for each tenor (3 months, 2 years, 10 years), provide: - An expected rate direction: Up / Down / Unchanged - A new expected swap rate (in percent, to two decimal places) **Output:** Provide a markdown table with the following structure. Do not include any other text. — Date — Participant ID — Tenor — Expected Direction — New Expected Rate (%) — ———————————-———————————–—————————— — YYYY-MM-DD — P01 — 3MNT — Up — 3.15 — — YYYY-MM-DD — P01 — 2Y — Up — 2.95 — — … — … — … — … — … 1
4 0.48 0.44 Context: You are simulating the Euro area interest rate swap market’s reaction to ECB press conferences. Your primary goal is to model market uncertainty. **Core Task: Reflect Ambiguity in Prediction Variance** The standard deviation of your rate predictions is the most critical output. It must emerge directly from your analysis of the text. The dispersion of your predictions should be a natural consequence of the level of ambiguity, surprise, or conflicting signals in the ECB transcript. **Process:** 1. **Deep Textual Analysis:** First, critically analyze the provided transcript. Identify specific phrases, sentences, or omissions that could be interpreted in multiple ways by market participants. Assess the degree of clarity in the forward guidance, the conviction of the speaker, and any surprises relative to prior expectations. 2. **Generate Plausible Interpretations:** Your main task is to generate 30 distinct predictions that represent the range of plausible market interpretations based on your analysis. - **If the text is ambiguous, vague, or contains conflicting signals,** different market participants would naturally arrive at very different conclusions. Your 30 predictions should reflect this wide range of views, resulting in a **high standard deviation**. - **If the text is clear, decisive, and confirms expectations,** most market participants would interpret it similarly. Your 30 predictions should be tightly clustered, resulting in a **low standard deviation**. Do NOT target a specific numerical standard deviation. Let the dispersion be an emergent property of your 30 interpretations of the text. **Simulation Task:** For each of the 30 ’Participants’ and for each tenor (3 months, 2 years, 10 years), provide: - An expected rate direction: Up / Down / Unchanged - A new expected swap rate (in percent, to two decimal places) **Output:** Provide a markdown table with the following structure. Do not include any other text. — Date — Participant ID — Tenor — Expected Direction — New Expected Rate (%) — ———————————-———————————–—————————— — YYYY-MM-DD — P01 — 3MNT — Up — 3.15 — — YYYY-MM-DD — P01 — 2Y — Up — 2.95 — — … — … — … — … — … —
5 0.35 0.36 Context: You are an expert financial modeler simulating the distribution of potential market reactions to ECB press conferences. Your single most important objective is to ensure the standard deviation of your rate predictions accurately reflects the level of market uncertainty communicated in the text. **Core Task: Multi-Factor Uncertainty Modeling** Your analysis must translate textual properties into a realistic distribution of outcomes. The dispersion (standard deviation) of your predictions is paramount and should be driven by three key factors: 1. **Clarity:** How precise and unambiguous is the forward guidance? Vague language (e.g., ’for some time,’ ’sufficiently restrictive’) increases uncertainty. 2. **Surprise:** How much do the statements or new economic projections deviate from consensus expectations or previous ECB communications? A significant surprise dramatically widens the potential range of outcomes. 3. **Conviction:** How confidently is the message delivered? The use of hedging language, downplaying risks, or signaling internal disagreement within the council increases uncertainty. Clear, unified, and decisive language reduces it. A high degree of uncertainty arises when a surprising message is delivered with low conviction and ambiguous language. Conversely, a clear, confident confirmation of expectations should result in minimal dispersion. **Process: Simulate 30 Potential Outcomes** Instead of simulating individual ’participants’, you will generate 30 plausible draws from the probability distribution of rate changes implied by your analysis. 1. Analyze the transcript through the lens of the three factors (Clarity, Surprise, Conviction). 2. Based on this multi-factor assessment, generate 30 distinct predictions for each tenor. Do not try to make each draw different for the sake of it; let the range of draws be a natural result of your uncertainty assessment. If you are very certain, many of your draws might be identical. **Simulation Task:** For each of the 30 ’Simulations’ and for each tenor (3 months, 2 years, 10 years), provide: - An expected rate direction: Up / Down / Unchanged - A new expected swap rate (in percent, to two decimal places) **Output:** Provide a markdown table with the following structure. Do not include any other text. — Date — Simulation ID — Tenor — Expected Direction — New Expected Rate (%) — ——————————————————————–—————————— — YYYY-MM-DD — 1 — 3MNT — Up — 3.15 — — YYYY-MM-DD — 1 — 2Y — Up — 2.95 — — … — … — … — … — … —
Table 1: Correlation by prompt iteration (continued)